Which normal distribution is also the standard normal curve?
The normal distribution would be a standard normal distribution if it had a mean of 0 and standard deviation of 1.
What are two events that cannot occur at the same time called?
Two events that cannot occur at the same time are called mutually exclusive.
When is data considered invalid?
Data is considered invalid when it is wrong or has changed. Data represents facts that are recorded, so any alterations made to it can make it invalid.
What does one standard deviation mean?
Standard deviation is a measure of variation from the mean of a data set. 1 standard deviation from the mean (which is usually + and - from mean) contains 68% of the data.
What is the probability of obtaining 4 ones in a row when rolling a fair six sided die?
The following is the probability of obtaining 4 ones IN THE FIRST FOUR rolls of a fair die.
Pr(4 1's) = Pr(1)*Pr(1)*Pr(1)*Pr(1) since the events are independent.
Pr(4 1's) = Pr(1)4 = (1/6)4 = 1/1296 = 0.000772
What is the number of distinguishable permutations of the letters in the word GLASSES?
The solution is count the number of letters in the word and divide by the number of permutations of the repeated letters; 7!/3! = 840.
If two events A and B are mutually exclusive what does the special rule of addition state?
If two events A and B are mutually exclusive, the special rule of addition states that the probability of one or the other event's occurring equals the sum of their probabilities. This rule is expressed in the following formula:
Special Rule of Addition
(5-2)
Equation 5-2
She can make 9! or 362880 lineups.
it depends on how many numbers are on the number cube. if it only goes to 6, then the probability is 1/6. if it goes to 8 then the probability is 2/8. and so on and so on...
How do you make a tree diagram to show the sample space after tossing four pennies?
You start with a point along the left edge of a page, halfway down. Form that point you have two line segments representing the outcomes of the first toss: a line going upwards (indicating Heads, H) and the other going downwards (indicating Tails, T). The letters H and T are written at the end of the line segments with the probability of each outcome written alongside the line. From the point marked H you repeat the two branches, representing the outcomes of the second toss and do the same from the T. From the two Hs and two Ts, repeat for the outcome of the third toss and then for the fourth. You should end up with a tree, going left to right, and ending with 16 (=2*2*2*2) outcomes.
It is a good idea to be systematics about drawing a tree diagram: always have the same outcome for the lines going up. That way you can be sure that you make mistakes.
What determines the z-score boundaries for the critical region?
Z-score boundaries are a part of the study of statistics. Z-scores are given by published tables. They refer to the proportion of values that lie between a number, Z and the mean.
How many different ways can you arrange the letters in the word fish?
There are 4 distinguishable letters in the word fish, so there is 4! or 24 different ways can you arrange the letters in the word fish.
If a 8 sided die is rolled 5 times how many different outcomes are there?
There are 85 or 32768 outcomes.
What is the probability of choosing a marble that is not blue in problem?
It can not be determined with the data provided.
Hector has a bag with 6 coins: 2 quarters, 1 nickel and 3 dimes.
The probability of randomly pulling a dime is: 3/6 = 1/2 = 0.50 = 50.0%
Hector's chance of pulling a quarter out of the bag is 2/6 = 1/3.
The probability that it is a spade when drawn is 13/52 or 1/4.
The probability of drawing an Ace from a 52 card deck is 4/52 or 1/13.